OpenStudy (anonymous):

use the rational root theorem then synthetic division (if necessary) to solve for all roots of the polynomial a) x^3-4x^2-4x+16 = f(x)

OpenStudy (anonymous):

x^3-4x^2-4x+16 x^2( x - 4) + 4( x + 4) = 0 (x + 4) ( x^2- 4) = 0 x = -4, x =2, x = =2

OpenStudy (anonymous):

and that is all you do?

OpenStudy (anonymous):

Done :)

OpenStudy (anonymous):

Thanks:)

OpenStudy (anonymous):

using the rational root theorem or p/q, you find that for -2,2,4 values for x you get zero. the rational root theorem is done by finding all that factors of 16/1 and inputing the factors back to the equation. the factors that gives you zero are the numbers that solve the problem.